OpenStudy (rocksamantha5):

Solve for x. 3(x + 2) + 4(x - 5) = 10

MsBrains (ms-brains):

Let's solve your equation step-by-step. 3(x+2)+4(x−5)=10 Step 1: Simplify both sides of the equation. 3(x+2)+4(x−5)=10 Simplify: (Show steps) 7x−14=10 Step 2: Add 14 to both sides. 7x−14+14=10+14 7x=24 Step 3: Divide both sides by 7. 7x/7=24/7 x=24/7

OpenStudy (rocksamantha5):

Thanks!